OpenStudy (anonymous):

Is the following statement true or false, and why? If f(x) =x^2-4/x-2 and g(x)=x+2, then f and g are the same function.

hero (hero):

Let me post a better solution.

hero (hero):

Assume f(x) = g(x) Let x = 2 Then f(2) = g(2) False Therefore f(x) ≠ g(x)

OpenStudy (anonymous):

Hold on, I think I got it. You're saying that I can plug in any number and if they don't have the same outcome they can't be the same. Is that it??

hero (hero):

Yes, correct.

hero (hero):

Because you can try any other number except x = 2 and it will work.

OpenStudy (anonymous):

My mind is blown at how easy that was, and you just came and... like nothing solved it. How would I answer the question though? Would I say "They are not the same because both functions have different outcomes"

OpenStudy (anonymous):

Or im thinking of putting, "f(2)=x^2-4/x-2 ≠ g(x)=x-2"

hero (hero):

No, the outcome would be the same for any other number except for x = 2. So they are the same with that one exception. f(x) has a discontinuity at x = 2. That's the only thing that makes them different. Another way to see it would be to graph both at the same time on your calculator.

hero (hero):

That one exception is the only thing that makes f(x) and g(x) different.

OpenStudy (anonymous):

Yeah, so by showing the exception I would be proving that they are not the same and answering the question. Correct?

hero (hero):

Correct

OpenStudy (anonymous):

Ok just to make sure now, "False, they are not the same because *show exception*"

OpenStudy (anonymous):

Thank you very much.